OpenStudy (anonymous):

solve the linear equations below:

OpenStudy (anonymous):

4x - 3y = 17 5x + 4y = 60

OpenStudy (ivancsc1996):

You just need to replace like this. First: 4x-3y=17 is 4x=17+3y is \[x=\frac{ 17 + 3y }{ 4 }\] The you insert that x into the second equation:\[(5 \times \frac{ 17 + 3y }{ 4 })+4y = 60 \rightarrow y= \frac{60-(5 \times \frac{ 17 + 3y }{ 4 }) }{ 4 }\]Then you solve for y and instert that back into the first equation and solve for x

OpenStudy (anonymous):

Thank you do much this is a very helpful way to solve linear equations :)